# Statistical Properties of a Polarization Vector’s Ellipticity Angle

### v1

@@ Abstract
The orientation of a polarization vector on the Poincaré sphere is defined by its position and ellipticity angles. The radio emission from pulsars, magnetars, and fast radio bursts can be elliptically polarized, and measurements of the ellipticity angle (EA) have become increasingly important in interpretations and models of their polarization. An in-depth understanding of the statistical properties of the measured polarization angles is a prerequisite to their detailed interpretation. While the statistics of the position angle (PA) have been understood for some time, the statistics of the EA do not appear to be as well developed as those of the PA. The statistical properties of the EA are derived when the amplitude of the polarization vector is constant, to include its probability density, mean, standard deviation, and confidence limits. Similarly to the PA, the standard deviation and confidence limits of the EA vary inversely with the polarization signal-to-noise ratio. However, unlike the PA, the probability density of the EA is generally asymmetric, its standard deviation and confidence limits are dependent upon the intrinsic value of the EA, and the measured EA is biased by the instrumental noise, particularly at low signal-to-noise ratios and large values of the intrinsic EA. General expressions for the joint probability density of the polarization angles and the probability density of the EA are also derived when the amplitude of the polarization vector fluctuates due to the superposition of incoherent modes of orthogonal polarization.
@@ 1 INTRODUCTION
Pulsar average profiles have typically shown how the total intensity, circular polarization, linear polarization, and polarization position angle (PA) vary across the pulsar’s pulse. In retrospect, measurements of the polarization vector’s ellipticity angle (EA) have been absent from most of these profiles. Both the PA and the EA are needed to completely define the orientation of the vector on the Poincaré sphere. The observations of PSR B0329+54 by Edwards & Stappers (2004) and PSR B0809+74 by Edwards (2004) were some of the first to investigate the behavior of the EA in detail. At some pulse longitudes in both pulsars, individual samples of the EA are concentrated near the equator of the Poincaré sphere, but the average EA makes a large excursion toward one of the poles of the sphere. These excursions tend to be accompanied by a large discontinuity in the PA of $\Delta\psi\simeq\pi/2$. The PA discontinuities are indicative of orthogonal modes of polarization (OPMs), which are generally attributed to the natural modes of wave propagation in the pulsar’s magnetosphere (Melrose & Stoneham 1977; Melrose 1979; Allen & Melrose 1982; Barnard & Arons 1986). The PA and EA in parts of PSR B0031-07 (Ilie et al., 2020) behave similarly to those of PSRs B0329+54 and B0809+74. The PA and EA in the central component of PSR B0329+54 combine to form an intriguing partial annulus in one hemisphere of the sphere. Edwards & Stappers (2004) suggested the annulus may be the result of generalized Faraday rotation in the pulsar magnetosphere (Kennett & Melrose 1998). They attributed the large EA excursions observed elsewhere in the pulse to either a transition between polarization modes that were not precisely orthogonal or an OPM transition combined with a circularly polarized emission component.
@@
More recent observations of pulsars, magnetars, and fast radio bursts (FRBs) have revealed large variations in the EA as functions of pulse longitude, frequency, or time. In some cases, the trajectories of their polarization vectors trace arcs or great circles on the Poincaré sphere. The magnitude of the EA can be large as the vector approaches a pole of the sphere. Dyks et al. (2021) attributed the longitude-dependent variations in the EA of PSR B1451-68 to changes in the relative intensity of coherent OPMs. They recognized that the variations in PA and EA caused by such an OPM transition follow a great circle on the Poincaré sphere, whereas variations due to changes in the phase offset between the modes generally trace a small circle on the sphere. Oswald et al. (2023) proposed that the OPMs in a pulsar’s magnetosphere are partially coherent and demonstrated that the frequency-dependent variations in the EA of PSR J0820-1350 and the longitude-dependent EA variations in PSR J1157-6224 are consistent with changes in the phase offset between the modes. They also demonstrated that the longitude-dependent variations in the EA of PSR J0134-2937 are consistent with a change in the relative intensity of the modes. Lower et al. (2024) found that frequency-dependent variations in the PA and EA of the magnetar XTE J1810-197 follow a circle on the Poincaré sphere. The possible explanations for this behavior include Faraday conversion in the relativistic, magnetized plasma that surrounds the magnetar or mode coupling within its magnetosphere. Bera et al. (2025) showed that temporal variations in the PA and EA of FRBs 20210912A and 20230708A follow great circles on the sphere. The variations are consistent with propagation effects in a birefringent medium located in the outer magnetosphere of the FRB progenitor or in the magnetized plasma surrounding it. Cao et al. (2025) observed large variations in the PA and EA across single pulses of PSR B1919+21. They attributed the variations to changes in the phase offset of partially coherent polarization modes. The aforementioned examples illustrate a variety of EA behaviors and emphasize the importance of the EA in the interpretation of the observations, particularly within the context of different polarization models.
@@
An in-depth understanding of the statistical properties of the polarization angles is a prerequisite to their detailed interpretation. The statistical properties of the PA have been understood for quite some time (e.g., Davenport & Root 1958; Papoulis 1965; Naghizadeh-Khouei & Clarke 1993). The probability density function (pdf) of the PA, $f_{\psi}(\psi)$, is unimodal and symmetric about its intrinsic value, $\psi_{o}$. When the signal-to-noise ratio (SNR) in linear polarization is large ($s\gg 1)$, $f_{\psi}(\psi)$ is Gaussian with a mean of $\psi_{o}$ and a standard deviation of $\sigma_{\psi}=1/(2s)$. The PA pdf is uniform when $s=0$. It is straightforward to numerically derive the confidence limits of $f_{\psi}(\psi)$, because it is symmetric about $\psi_{o}$ (Wardle & Kronberg 1974; Naghizadeh-Khouei & Clarke 1993; Everett & Weisberg 2001). And since the pdf is Gaussian at high SNR, the confidence limits can be approximated by its standard deviation. McKinnon & Stinebring (1998) generalized $f_{\psi}(\psi)$ to include OPMs. This pdf is generally bimodal. The angular separation between the two peaks of the pdf is $\Delta\psi=\pi/2$, the difference between the heights of the peaks is proportional to the difference in mean intensities of the modes, and the widths of the peaks vary inversely with the polarization fluctuations (McKinnon 2022).
@@
The purpose of this paper is to derive the general statistical properties of the EA, to include its pdf, mean, standard deviation, and confidence limits. The paper is organized as follows. A general expression for the joint probability density of the PA and EA is derived in Section 2 for the case when the amplitude of the polarization vector is constant. The pdf of the EA is derived from the joint density. The mean and standard deviation of the EA are then calculated from its pdf. The EA statistics derived in this section are complementary to the PA statistics documented in Naghizadeh-Khouei & Clarke (1993). In Section 3, the joint density and EA pdf are derived for the case when the amplitude of the polarization vector fluctuates due to the superposition of incoherent OPMs. The EA statistics developed in this section are complementary to those of the PA derived for OPMs in McKinnon & Stinebring (1998). The results of the analysis are compared with observations in Section 4. The statistical properties of the PA and EA are compared and contrasted in Section 5. Summary comments are listed in Section 6. A method for calculating the confidence limits of the EA is proposed in the Appendix.
2 Ellipticity Angle of a Fixed Polarization Vector
@@ 2.1 Joint Probability Density of the Polarization Angles
The procedure for deriving the joint probability density of the polarization angles is outlined in McKinnon (2003). To briefly summarize that analysis, the Stokes parameters $Q$, $U$, and $V$ are independent, Gaussian random variables when the amplitude and orientation of the polarization vector are fixed. Their joint probability density is then the product of their individual distributions. The joint probability density of the polarization angles can be found by converting the Cartesian coordinates of the Stokes parameters to spherical coordinates and integrating over radius. For a polarization vector with amplitude $\mu_{p}$, an intrinsic PA $\psi_{o}$, and an intrinsic EA $\chi_{o}$, the general expression for the joint probability density of $\psi$ and $\chi$ is
@@
2.2 Probability Density of the Ellipticity Angle
Figure 2: Examples of the EA probability density function (pdf) when the amplitude of the polarization vector is constant. The top panel shows the pdf for different values of the intrinsic EA, $\chi_{o}$, at a constant value of the signal-to-noise ratio in total polarization, $s=\mu_{p}/\sigma_{n}=4$. The bottom panel shows the pdf for different values of $s$ when $\chi_{o}$ is held constant at $30^{\circ}$.
The pdf of $\chi$, $f_{\chi}(\chi)$, can be derived by integrating Equation 1 over $\psi$. The resulting pdf is a function of $s$ and $\chi_{o}$. The pdfs for $\chi_{o}=0$ and $\chi_{o}=\pi/4$ can be derived analytically. The pdfs for other values of $\chi_{o}$ can be calculated numerically. Examples of $f_{\chi}(\chi)$ are illustrated in Figure 2. The top panel of the figure shows how $f_{\chi}(\chi)$ evolves as the intrinsic EA changes from $\chi_{o}=0$ to $\chi_{o}=\pi/4$ at a fixed value of SNR in total polarization. The pdfs are symmetric and Gaussian-like for small to intermediate values of $\chi_{o}$. The pdfs become slightly asymmetric as $\chi_{o}$ increases and evolve to what resembles a mirror image of a Rayleigh distribution at $\chi_{o}=\pi/4$. The bottom panel of the figure demonstrates how the pdf evolves with increasing SNR while the intrinsic EA remains constant at $\chi_{o}=\pi/6$. The low to moderate values of $s$ used in the panel were chosen to show that the shape of the pdf changes significantly in this SNR regime. The pdf is a cosine distribution when the radiation is unpolarized ($s=0$). As the SNR increases, the pdf is initially asymmetric and becomes increasingly symmetric until it resembles a Gaussian distribution at high SNR.
@@ 2.2 Probability Density of the Ellipticity Angle
The standard deviation of $f_{\theta}(\theta)$ is $\sigma_{\theta}=s^{-1}\sqrt{2-\pi/2}$. The standard deviation of $\chi$ is one-half this value:
@@ 2.3 Measured Ellipticity Angle
Examples of $\chi_{m}$ calculated from Equation 12 are shown in Figure 3. The dashed green line in the top panel of the figure shows the dependence of the measured EA upon $s$ when $\chi_{o}=15^{\circ}$. The bottom panel compares the measured EA to the intrinsic EA for different values of $s$. The measured value of the EA systematically underestimates the intrinsic value, particularly at low SNR in total polarization and at large values of $\chi_{o}$. The difference between the intrinsic and measured values of the EA is caused by a contribution of the instrumental noise to the mean linear polarization. This problem is well known, and a variety of methods have been developed to reduce the effect of this instrumental bias (e.g., Simmons & Stewart 1985; Everett & Weisberg 2001; Plaszczynski et al., 2015). The method proposed by Everett & Weisberg (2001), hereafter the EW estimator, is commonly used to compensate the measured linear polarization ($L_{m}=\langle L\rangle$) of pulsars for instrumental noise. Their estimate of the true linear polarization, $L_{t}$, from $L_{m}$ is given by
@@
The solid red line in the top panel of Figure 3 shows the corrected value of $\chi_{m}$ after Equation 13 has been applied to the linear polarization measurement. The corrected value slightly underestimates the actual value of $\chi_{o}$ at low SNR, but converges to $\chi_{o}$ at $s>3$. The corrected value of $\chi_{m}$ is not shown for $s<1.81$, because Equation 13 sets $L_{t}=0$ for $s\cos(2\chi_{o})<1.57$. Plaszczynski et al. (2015) proposed a modified asymptotic (MAS) estimator to compensate the measured linear polarization for instrumental noise. The MAS estimator calculates $L_{t}$ from $L_{m}$ using
@@
2.4 Standard Deviation of the Ellipticity Angle’s Probability Density
@@
Approximations to $\sigma_{\chi}$ can be determined from the definition of $\chi$ (Equation 10) using the standard propagation of errors (e.g., Naghizadeh-Khouei & Clarke 1993; Cao et al., 2025):
@@
For circularly polarized radiation ($\chi_{o}=\pi/4$), $\sigma_{l}=\sigma_{n}\sqrt{2-\pi/2}$, and the approximation calculated from Equation 16 is the same result given by Equation 9. The approximation is shown by the open circles in Figure 4. It reproduces the actual standard deviation for SNRs of $s>4$.
@@
3 Ellipticity Angle from Polarization Fluctuations
@@ 3.1 OPM Statistical Model
Here, the random variables $X_{N,Q}$, $X_{N,U}$, and $X_{N,V}$ represent the independent instrumental noise in each of the measured Stokes parameters $Q$, $U$, and $V$. The magnitude of these random variables is assumed to be equal with a value of $\sigma_{n}$. The equations are written assuming mode A is the stronger of the two modes. Therefore, the PA of mode A is $\psi_{o}$, and its EA is $\chi_{o}$. The PA of mode B is $\psi_{o}\pm\pi/2$ and its EA is $\chi_{o}$. The fluctuations in mode intensities cause the amplitude of the resultant polarization vector to fluctuate. Since the OPMs are assumed to be superposed and incoherent, the average SNR in total polarization is determined by the difference in mode mean intensities. Therefore, while the SNR of a single polarization sample can be very large, the SNR on average will be small when the mode mean intensities are comparable.

### v2

@@ Abstract
The orientation of a polarization vector on the Poincaré sphere is defined by its position angle (PA) and ellipticity angle (EA). The radio emission from pulsars, magnetars, and fast radio bursts can be elliptically polarized, and measurements of the EA have become increasingly important in interpretations and models of their polarization. An in-depth understanding of the statistical properties of the measured polarization angles is a prerequisite to their detailed interpretation. While the statistics of the PA have been understood for some time, the statistics of the EA do not appear to be as well developed as those of the PA. The statistical properties of the EA are derived when the amplitude of the polarization vector is constant, to include its probability density, mean, standard deviation, and confidence limits. Similar to the PA, the standard deviation and confidence limits of the EA vary inversely with the polarization signal-to-noise ratio. However, unlike the PA, the probability density of the EA is generally asymmetric, its standard deviation and confidence limits are dependent upon the intrinsic value of the EA, and the measured EA is biased by the instrumental noise, particularly at low signal-to-noise ratios and large values of the intrinsic EA. General expressions for the joint probability density of the polarization angles and the probability density of the EA are also derived when the amplitude of the polarization vector fluctuates due to the superposition of incoherent modes of orthogonal polarization.
@@ 1 INTRODUCTION
Pulsar average profiles have typically shown how the total intensity, circular polarization, linear polarization, and polarization position angle (PA) vary across the pulsar’s pulse. In retrospect, measurements of the polarization vector’s ellipticity angle (EA) have been absent from most of these profiles. Both the PA and the EA are needed to completely define the orientation of the vector on the Poincaré sphere. The observations of PSR B0329+54 by Edwards & Stappers (2004) and PSR B0809+74 by Edwards (2004) were some of the first to investigate the behavior of the EA in detail. At some pulse longitudes in both pulsars, individual samples of the EA are concentrated near the equator of the Poincaré sphere, but the average EA makes a large excursion toward one of the poles of the sphere. These excursions tend to be accompanied by a large discontinuity in the PA of $\Delta\psi\simeq\pi/2$. The PA discontinuities are indicative of orthogonal modes of polarization (OPMs), which are generally attributed to the natural modes of wave propagation in the pulsar’s magnetosphere (Melrose & Stoneham 1977; Melrose 1979; Allen & Melrose 1982; Barnard & Arons 1986). The PA and EA in parts of PSR B0031-07 (Ilie et al. 2020) behave similarly to those of PSRs B0329+54 and B0809+74. The PA and EA in the central component of PSR B0329+54 combine to form an intriguing partial annulus in one hemisphere of the sphere. Edwards & Stappers (2004) suggested the annulus may be the result of generalized Faraday rotation in the pulsar magnetosphere (Kennett & Melrose 1998). They attributed the large EA excursions observed elsewhere in the pulse to either a transition between polarization modes that were not precisely orthogonal or an OPM transition combined with a circularly polarized emission component.
@@
More recent observations of pulsars, magnetars, and fast radio bursts (FRBs) have revealed large variations in the EA as functions of pulse longitude, frequency, or time. In some cases, the trajectories of their polarization vectors trace arcs or great circles on the Poincaré sphere. The magnitude of the EA can be large as the vector approaches a pole of the sphere. Dyks et al. (2021) attributed the longitude-dependent variations in the EA of PSR B1451-68 to changes in the relative intensity of coherent OPMs. They recognized that the variations in PA and EA caused by such an OPM transition follow a great circle on the Poincaré sphere, whereas variations due to changes in the phase offset between the modes generally trace a small circle on the sphere. Oswald et al. (2023) proposed that the OPMs in a pulsar’s magnetosphere are partially coherent and demonstrated that the frequency-dependent variations in the EA of PSR J0820-1350 and the longitude-dependent EA variations in PSR J1157-6224 are consistent with changes in the phase offset between the modes. They also demonstrated that the longitude-dependent variations in the EA of PSR J0134-2937 are consistent with a change in the relative intensity of the modes. Lower et al. (2024) found that frequency-dependent variations in the PA and EA of the magnetar XTE J1810-197 follow a circle on the Poincaré sphere. The possible explanations for this behavior include Faraday conversion in the relativistic magnetized plasma that surrounds the magnetar or mode coupling within its magnetosphere. Bera et al. (2025) showed that temporal variations in the PA and EA of FRBs 20210912A and 20230708A follow great circles on the sphere. The variations are consistent with propagation effects in a birefringent medium located in the outer magnetosphere of the FRB progenitor or in the magnetized plasma surrounding it. Cao et al. (2025) observed large variations in the PA and EA across single pulses of PSR B1919+21. They attributed the variations to changes in the phase offset of partially coherent polarization modes. The aforementioned examples illustrate a variety of EA behaviors and emphasize the importance of the EA in the interpretation of the observations, particularly within the context of different polarization models.
@@
An in-depth understanding of the statistical properties of the polarization angles is a prerequisite to their detailed interpretation. The statistical properties of the PA have been understood for quite some time (e.g., Davenport & Root 1958; Papoulis 1965; Naghizadeh-Khouei & Clarke 1993). The probability density function (pdf) of the PA, $f_{\psi}(\psi)$, is unimodal and symmetric about its intrinsic value, $\psi_{o}$. When the signal-to-noise ratio (SNR) in linear polarization is large ($s\gg 1)$, $f_{\psi}(\psi)$ is Gaussian, with a mean of $\psi_{o}$ and a standard deviation of $\sigma_{\psi}=1/(2s)$. The PA pdf is uniform when $s=0$. It is straightforward to numerically derive the confidence limits of $f_{\psi}(\psi)$, because it is symmetric about $\psi_{o}$ (Wardle & Kronberg 1974; Naghizadeh-Khouei & Clarke 1993; Everett & Weisberg 2001). And since the pdf is Gaussian at high SNR, the confidence limits can be approximated by its standard deviation. McKinnon & Stinebring (1998) generalized $f_{\psi}(\psi)$ to include OPMs. This pdf is generally bimodal. The angular separation between the two peaks of the pdf is $\Delta\psi=\pi/2$, the difference between the heights of the peaks is proportional to the difference in the mean intensities of the modes, and the widths of the peaks vary inversely with the polarization fluctuations (McKinnon 2022).
@@
The purpose of this paper is to derive the general statistical properties of the EA, to include its pdf, mean, standard deviation, and confidence limits. The paper is organized as follows. A general expression for the joint probability density of the PA and EA is derived in Section 2 for the case when the amplitude of the polarization vector is constant. The pdf of the EA is derived from the joint density. The mean and standard deviation of the EA are then calculated from its pdf. The EA statistics derived in this section are complementary to the PA statistics documented in Naghizadeh-Khouei & Clarke (1993). In Section 3, the joint density and EA pdfs are derived for the case when the amplitude of the polarization vector fluctuates due to the superposition of incoherent OPMs. The EA statistics developed in this section are complementary to those of the PA derived for OPMs in McKinnon & Stinebring (1998). The results of the analysis are compared with observations in Section 4. The statistical properties of the PA and EA are compared and contrasted in Section 5. Summary comments are listed in Section 6. A method for calculating the confidence limits of the EA is proposed in the Appendix.
2 EA of a Fixed Polarization Vector
@@ 2.1 Joint Probability Density of the Polarization Angles
The procedure for deriving the joint probability density of the polarization angles is outlined in McKinnon (2003). To briefly summarize that analysis, the Stokes parameters $Q$, $U$, and $V$ are independent Gaussian random variables when the amplitude and orientation of the polarization vector are fixed. Their joint probability density is then the product of their individual distributions. The joint probability density of the polarization angles can be found by converting the Cartesian coordinates of the Stokes parameters to spherical coordinates and integrating over radius. For a polarization vector with amplitude $\mu_{p}$, an intrinsic PA $\psi_{o}$, and an intrinsic EA $\chi_{o}$, the general expression for the joint probability density of $\psi$ and $\chi$ is
@@
2.2 Probability Density of the EA
Figure 2: Examples of the EA pdf when the amplitude of the polarization vector is constant. The top panel shows the pdfs for different values of the intrinsic EA, $\chi_{o}$, at a constant value of the SNR in total polarization, $s=\mu_{p}/\sigma_{n}=4$. The bottom panel shows the pdfs for different values of $s$ when $\chi_{o}$ is held constant at $30^{\circ}$.
The pdf of $\chi$, $f_{\chi}(\chi)$, can be derived by integrating Equation 1 over $\psi$. The resulting pdf is a function of $s$ and $\chi_{o}$. The pdfs for $\chi_{o}=0$ and $\chi_{o}=\pi/4$ can be derived analytically. The pdfs for other values of $\chi_{o}$ can be calculated numerically. Examples of $f_{\chi}(\chi)$ are illustrated in Figure 2. The top panel of the figure shows how $f_{\chi}(\chi)$ evolves as the intrinsic EA changes from $\chi_{o}=0$ to $\chi_{o}=\pi/4$ at a fixed value of SNR in total polarization. The pdfs are symmetric and Gaussian-like for small to intermediate values of $\chi_{o}$. The pdfs become slightly asymmetric as $\chi_{o}$ increases and evolve to what resembles a mirror image of a Rayleigh distribution at $\chi_{o}=\pi/4$. The bottom panel of the figure demonstrates how the pdf evolves with increasing SNR while the intrinsic EA remains constant at $\chi_{o}=\pi/6$. The low to moderate values of $s$ used in the panel were chosen to show that the shape of the pdf changes significantly in this SNR regime. The pdf is a cosine distribution when the radiation is unpolarized ($s=0$). As the SNR increases, the pdf is initially asymmetric and becomes increasingly symmetric, until it resembles a Gaussian distribution at high SNR.
@@ 2.2 Probability Density of the EA
The standard deviation of $f_{\theta}(\theta)$ is $\sigma_{\theta}=s^{-1}\sqrt{2-\pi/2}$. The standard deviation of $\chi$ is one-half of this value:
@@ 2.3 Measured EA
Examples of $\chi_{m}$ calculated from Equation 12 are shown in Figure 3. The dashed green line in the top panel of the figure shows the dependence of the measured EA upon $s$ when $\chi_{o}=15^{\circ}$. The bottom panel compares the measured EA to the intrinsic EA for different values of $s$. The measured value of the EA systematically underestimates the intrinsic value, particularly at low SNR in total polarization and at large values of $\chi_{o}$. The difference between the intrinsic and measured values of the EA is caused by a contribution of the instrumental noise to the mean linear polarization. This problem is well known, and a variety of methods have been developed to reduce the effect of this instrumental bias (e.g., Simmons & Stewart 1985; Everett & Weisberg 2001; Plaszczynski et al., 2014). The method proposed by Everett & Weisberg (2001), hereafter the EW estimator, is commonly used to compensate the measured linear polarization ($L_{m}=\langle L\rangle$) of pulsars for instrumental noise. Their estimate of the true linear polarization, $L_{t}$, from $L_{m}$ is given by
@@
The solid red line in the top panel of Figure 3 shows the corrected value of $\chi_{m}$ after Equation 13 has been applied to the linear polarization measurement. The corrected value slightly underestimates the actual value of $\chi_{o}$ at low SNR but converges to $\chi_{o}$ at $s>3$. The corrected value of $\chi_{m}$ is not shown for $s<1.81$, because Equation 13 sets $L_{t}=0$ for $s\cos(2\chi_{o})<1.57$. Plaszczynski et al. (2014) proposed a modified asymptotic (MAS) estimator to compensate the measured linear polarization for instrumental noise. The MAS estimator calculates $L_{t}$ from $L_{m}$ using
@@
2.4 Standard Deviation of the EA’s Probability Density
@@
Approximations to $\sigma_{\chi}$ can be determined from the definition of $\chi$ (Equation 10) using the standard propagation of errors (e.g., Naghizadeh-Khouei & Clarke 1993; Cao et al. 2025):
@@
For circularly polarized radiation ($\chi_{o}=\pi/4$), $\sigma_{l}=\sigma_{n}\sqrt{2-\pi/2}$, and the approximation calculated from Equation 16 is the same result as given by Equation 9. The approximation is shown by the open circles in Figure 4. It reproduces the actual standard deviation for SNRs of $s>4$.
@@
3 EA from Polarization Fluctuations
@@ 3.1 OPM Statistical Model
Here, the random variables $X_{N,Q}$, $X_{N,U}$, and $X_{N,V}$ represent the independent instrumental noise in each of the measured Stokes parameters $Q$, $U$, and $V$. The magnitude of these random variables is assumed to be equal, with a value of $\sigma_{n}$. The equations are written assuming mode A is the stronger of the two modes. Therefore, the PA of mode A is $\psi_{o}$, and its EA is $\chi_{o}$. The PA of mode B is $\psi_{o}\pm\pi/2$ and its EA is $\chi_{o}$. The fluctuations in mode intensities cause the amplitude of the resultant polarization vector to fluctuate. Since the OPMs are assumed to be superposed and incoherent, the average SNR in total polarization is determined by the difference in the mode mean intensities. Therefore, while the SNR of a single polarization sample can be very large, the SNR on average will be small when the mode mean intensities are comparable.